A given task was estimated to cost $15,750 to complete. Currently, the project team has completed approximately 65% of the task, and have spent $10,500. a. Determine the indices: AC, PV, and EV b. Derive the values for cost and schedule variances, as well as cost performance and schedule performance indices. Is this project ahead or behind on schedule, and is it ahead or behind on budget? c. How much additional money will be required to finish this project? d. Estimate the total FINAL cost of completing this project (EAC). Is this within acceptable limits? Explain
Answer #1

a)

PV = $ 15,750

EV = 15750*0.65 = $ 10,237.5

AC = $ 10,500

b)

Cost Variance, CV = EV - AC

= 10237.5 - 10500

= $ -262.5

Schedule Variance, SV = EV - PV

= 10237.5 - 15750

= $ -5,512.5

Cost Variance and Schedule Variance are both negative. Therefore, the project is both behind schedule and over (behind) budget.

c)

Cost Performance Index, CPI = EV/AC = 10237.5/10500 = 0.975

Additional money required to finish this project, ETC = (BAC-EV)/CPI   (in this case, BAC = PV = 15750)

= (15750-10500)/0.975

= $ 5,385

d) EAC = AC+ETC

= 10500+5385

= $ 15,885